# FIAN/TD/04–25

### v1

@@ 1 Introduction
The aim of this letter is to point out that there exists an object (field) very different from the fields usually used in the Standard Model (SM), that may, in principle, affect general picture of the origin of some of phenomena in high-energy physics, cosmology and astrophysics. The main actor of this letter is the so-called singleton discovered by Dirac in 1963 in [1] [^1] The name was given even earlier in the math literature [4] to emphasize its simplicity from the representation theory perspective. as a specific branch of the solutions of a certain wave equation, that survives at infinity of $AdS_{4}$. There are two singleton fields of boson and fermion types often called $Rac$ and $Di$, respectively [5] (see [6, 7] for reviews). Later on it was realized that $Rac$ and $Di$ are nothing else as, respectively, free massless scalar and spinor fields in the 2+1 dimensional space [^2] See, e.g., [8]-[12] where singleton supermultiplets were considered in the context of supermembrane and higher-spin theory. presently identified with the boundary of $AdS_{4}$ within the paradigm of holographic correspondence [13]-[15], which interpretation is particularly relevant in the context of higher-spin holography [16, 17].
@@
Firstly, we interpret singletons as relativistic fields in $(A)dS_{d+1}$ providing for them Lorentz covariant field equations and actions which, to the best of our knowledge, were not available before. This is achieved within the unfolded dynamics approach (see [18] for an elementary review and more references). The proposed formulation differs from the Flato-Fronsdal dipole one [2] given in terms of a certain fourth-order field equation in $AdS_{d+1}$. In particular, it is free from the huge gauge symmetry of [2] eliminating local degrees of freedom in $AdS_{d+1}$.
@@ 2 Unfolded dynamics
The idea of the unfolded formulation was put forward in [23] where it was realized that the full system of nonlinear equations for massless HS gauge fields can be searched in the form (2.2) as a deformation of the free unfolded equations.
@@
where the gauge parameter $\varepsilon^{\Omega}(x)$ is a $(p_{\Omega}-1)$-form if $W^{\Omega}$ was a $p_{\Omega}$-form. Strictly speaking, this is true for the class of universal unfolded systems in which the compatibility conditions (2.3) hold independently of the dimension $d$ of space-time, i.e.,  (2.3) should be true disregarding the fact that any $(d+1)$-form is zero. This is the case in all unfolded systems considered in the HS literature including this paper.
@@ 8 Singleton as a beyond SM actor
As such, it is a sum of the three-form Lagrangians $L^{Rac}$ and $L^{Di}$ and four-form Lagrangians for the genuine $4d$ local fields contributing to $L^{SM}$ and $L^{GR}$. Let us stress that since the Lagrangians are closed, the respective actions are insensitive to the local variations of the three-cycles over which $L^{Rac}$ and $L^{Di}$ are integrated in the action,
@@
That bilinears of singletons contain graviton and an inert massless scalar may have observable consequences via, e.g., an additionally induced gravitational field as well as invisible (dark) matter. In other words, though the $4d$ singleton fields cannot be localized themselves in some region of a galaxy, their nonlinear combinations can induce localizable fields. Of course, to evaluate such effects one has to introduce and analyse interactions. The first step towards gravitational interaction consists of the covariantization of $D_{x}$ in (7.7) beyond the flat $AdS_{d+1}$ connection. Note, however, that, as usual in the unfolded dynamics, this is only the first step that demands further extension to respect the formal consistency in the sense of (2.3) beyond the linearized approximation. The simplest way towards solution to this problem is probably via construction of the unfolded version of $3d$ conformal gravity theory with $3d$ massless matter. The $4d$ singleton theory then will result via the application of the space-time metamorphoses mechanism of Section 3 to the $3d$ conformal theory. For some progress in this direction within $3d$ conformal higher-spin gravity see [43, 44], while unfolded formulation of conformal geometry was considered in [45]. The problem is to evaluate the effect of condensation of the gravitational and scalar fields in the galaxy induced by the singleton.
@@
The presence of complex coefficients in the singleton equations may be related to such a persisting problem in cosmology and high-energy physics as baryon asymmetry. Indeed, these properties fit the prominent Sakharov necessary conditions [48] for baryon asymmetry (for review see e.g. [49]). Positive cosmological constant may provide a non-equilibrium regime. Moreover, singletons endowed with appropriate inner structure may induce violation of the baryon number conservation. The presence of complex coefficients in the Lagrangian may induce $CP$ violation. Though, because of its $4d$ nonlocality, the singleton field can hardly be seen in usual scattering phenomena, it can affect the formation of a charged matter via a nonlinear manifestation of the Flato-Fronsdal theorem.
@@ Acknowledgments
I wish to thank Konstantin Alkalaev, Mikhail Alfimov, Andrei Barvinsky, Olga Gelfond, Sergey Godunov, Andrei Mironov, Nikita Misuna, Roman Nevzorov, Dmitry Ponomarev, Aleksandr Pukhov, Arkady Tseytlin, Boris Voronov, and Mikhail Vysotsky for useful comments. I am grateful for hospitality to Ofer Aharony, Theoretical High Energy Physics Group of Weizmann Institute of Science where a part of this work has been done. This work was supported by Theoretical Physics and Mathematics Advancement Foundation BASIS Grant No 24-1-1-9-5.

### v2

@@ 1 Introduction
The aim of this letter is to point out that there exists an object (field) very different from the fields usually used in the Standard Model (SM), that may, in principle, affect general picture of the origin of some of phenomena in high-energy physics, cosmology and astrophysics. The main actor of this letter is the so-called singleton discovered by Dirac in 1963 in [1] [^1] The name was given even earlier in the math literature [4] to emphasize its simplicity from the representation theory perspective. as a specific branch of the solutions of a certain wave equation, that survives at infinity of $AdS_{4}$. There are two singleton fields of boson and fermion types often called $Rac$ and $Di$, respectively [5] (see [6, 7] for reviews). Later on it was realized that $Rac$ and $Di$ and their supermultiplets considered in the context of supergravity $S^{7}$ compactifications, supermembrane and higher-spin theory, for instance, in [8]-[15] are nothing else as, respectively, free massless scalar and spinor fields in the 2+1 dimensional space presently identified with the boundary of $AdS_{4}$ within the paradigm of holographic correspondence [16]-[18], which interpretation is particularly relevant in the context of higher-spin holography [19, 20].
@@
Firstly, we interpret singletons as relativistic fields in $(A)dS_{d+1}$ providing for them Lorentz covariant field equations and actions which, to the best of our knowledge, were not available before. This is achieved within the unfolded dynamics approach (see [21] for an elementary review and more references). The proposed formulation differs from the Flato-Fronsdal dipole one [2] given in terms of a certain fourth-order field equation in $AdS_{d+1}$. In particular, it is free from the huge gauge symmetry of [2] eliminating local degrees of freedom in $AdS_{d+1}$. On the other hand, kinematically our formulation has some similarities with the interpretation of holography proposed in the framework of the nonlinear realisation approach in [22, 23] (though not at the Lagrangian level).
@@ 2 Unfolded dynamics
The idea of the unfolded formulation was put forward in [28] where it was realized that the full system of nonlinear equations for massless higher-spin gauge fields can be searched in the form (2.2) as a deformation of the free unfolded equations.
@@
where the gauge parameter $\varepsilon^{\Omega}(x)$ is a $(p_{\Omega}-1)$-form if $W^{\Omega}$ was a $p_{\Omega}$-form. Strictly speaking, this is true for the class of universal unfolded systems in which the compatibility conditions (2.3) hold independently of the dimension $d$ of space-time, i.e.,  (2.3) should be true disregarding the fact that any $(d+1)$-form is zero. This is the case in all unfolded systems considered in the higher-spin literature including this paper.
@@ 8 Singleton as a beyond SM actor
As such, it is a sum of the three-form Lagrangians $L^{Rac}$ and $L^{Di}$ and four-form Lagrangians for the genuine $4d$ local fields contributing to $L^{SM}$ and $L^{GR}$. Let us stress that since the Lagrangians are closed forms, the respective actions are insensitive to the local variations of the three-cycles over which $L^{Rac}$ and $L^{Di}$ are integrated in the action,
@@
That bilinears of singletons contain graviton and an inert massless scalar may have observable consequences via, e.g., an additionally induced gravitational field as well as invisible (dark) matter. In other words, though the $4d$ singleton fields cannot be localized themselves in some region of a galaxy, their nonlinear combinations can induce localizable fields. Of course, to evaluate such effects one has to introduce and analyse interactions. The first step towards gravitational interaction consists of the covariantization of $D_{x}$ in (7.66) beyond the flat $AdS_{d+1}$ connection. Note, however, that, as usual in the unfolded dynamics, this is only the first step that demands further extension to respect the formal consistency in the sense of (2.3) beyond the linearized approximation. The simplest way towards solution to this problem is probably via construction of the unfolded version of $3d$ conformal gravity theory with $3d$ massless matter. The $4d$ singleton theory then will result via the application of the space-time metamorphoses mechanism of Section 3 to the $3d$ conformal theory. For some progress in this direction within $3d$ conformal higher-spin gravity see [48, 49], while unfolded formulation of conformal geometry was considered in [50]. The construction of the nonlinear $3d$ conformal higher-spin theory interacting with singletons was proposed in the recent papers [51, 52]. The problem is to evaluate the effect of condensation of the gravitational and scalar fields in the galaxy induced by the singleton.
An intriguing group-theoretic singleton phenomenon found in [14] is that the compactification of eleven-dimensional supergravity on the seven sphere leads to singleton representations along with the usual $4d$ relativistic representations in the spectrum. Moreover, in the squashed seven sphere case singletons were argued to be involved into certain higgsing together with the usual $4d$ unitary representations. Since the field-theoretic interpretation of this phenomenon is still not clear it would be interesting to apply the developed unfolded machinery to its further analysis. The formalism developed in this paper may be most appropriate in this respect since unfolded dynamics essentially maps the representation theory to dynamical equations. Also a potentially useful for better understanding of this phenomenon is the coset space approach to $AdS/CFT$ of [23]. It would be interesting to elaborate more on its relation to the unfolded dynamics approach used in this paper.
@@
The presence of complex coefficients in the singleton equations may be related to such a persisting problem in cosmology and high-energy physics as baryon asymmetry. Indeed, these properties fit the prominent Sakharov necessary conditions [55] for baryon asymmetry (for review see, e.g., [56]). Positive cosmological constant may provide a non-equilibrium regime. Moreover, singletons endowed with appropriate inner structure may induce violation of the baryon number conservation. The presence of complex coefficients in the Lagrangian may induce $CP$ violation. Though, because of its $4d$ nonlocality, the singleton field can hardly be seen in usual scattering phenomena, it can affect the formation of a charged matter via a nonlinear manifestation of the Flato-Fronsdal theorem.
@@ Acknowledgments
I wish to thank Konstantin Alkalaev, Mikhail Alfimov, Andrei Barvinsky, Olga Gelfond, Sergey Godunov, Carlo Iazeolla, Evgeny Ivanov, Andrei Mironov, Nikita Misuna, Roman Nevzorov, Bengt Nilsson, Dmitry Ponomarev, Aleksandr Pukhov, Arkady Tseytlin, Boris Voronov, and Mikhail Vysotsky for useful comments. I am grateful for hospitality to Ofer Aharony, Theoretical High Energy Physics Group of Weizmann Institute of Science where a part of this work has been done. This work was supported by Theoretical Physics and Mathematics Advancement Foundation BASIS Grant No 24-1-1-9-5.